OpenStudy (anonymous):

What are the asymptotes of (x^2/144)-(y^2/4)=1? I think that it is y=1/6x? Am I on the right track?

OpenStudy (anonymous):

I think there are two asymptotes !

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

12y=+or-2x

OpenStudy (anonymous):

I dont understand...

OpenStudy (anonymous):

I understand the ± but not the 12y and the 2x.

OpenStudy (anonymous):

it's an equation of a hyperbola It should be "written" in this form \[(x/a)^2-(y/b)^2=\pm 1\] then the asymptotes are \[ay=\pm bx\]

OpenStudy (anonymous):

\[x^2/144=(x/12)^2\]

OpenStudy (anonymous):

Did you understand now ?